OpenStudy (anonymous):

Solve 2cos²x-1-cosx=0 Interval= 0≤x<2π Please show steps

OpenStudy (anonymous):

let\[t=\cos x\]u have\[2t^2-t-1=0\]now factor it or solve by quadratic equation formula

OpenStudy (anonymous):

im supposed to end up with radians for my answer

OpenStudy (anonymous):

@Shi

OpenStudy (anonymous):

What mukushla said is a good step to get your answer, which will be in radians. Can you factor that equation?

OpenStudy (anonymous):

Yea I just factored it. Now im stuck finding the solutions within the interval.

OpenStudy (anonymous):

ok what did you factor it into? you dont have to worry about the interval just yet

OpenStudy (anonymous):

x=1/2 x=1

OpenStudy (anonymous):

you basically got it, I think the correct answers up to that point are t = -1/2 and t = 1 I used t just so we dont get confused ok after that you convert t back into cos(x)

OpenStudy (anonymous):

so you have cos(x) = -1/2 and cos(x) = 1 following so far?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we look at cos(x) = 1 first at what value of x does cos(x) = 1?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

yup, thats one of your answers. same thing for the second equation, cos(x) = -1/2 I recommend looking at a unit circle since there are 2 values, and taking the inverse of both sides only gives you one answer

OpenStudy (anonymous):

Ok thanks, i figured it out. Its not as bad as it seems lol

OpenStudy (anonymous):

haha yeah, figuring out what to do is always the hard part. gl with the rest of your hw

OpenStudy (anonymous):

Thanks